VT009

PI: Gurvits
Title: VSOP-test

Session: Feb 1997 (session 1)
Frequency: 6 cm
Mode:
Obs time (UT): EU 09:30(20/02)-12:00(20/02)
Stations: EbJbMcNtOnShTrWb
Processor: Soc
EU Mk3 Tapes/St.: 1

Comment


To give feedback on this experiment, click here

Experiment Feedback

Effelsberg (dgraham at Thu Feb 20 07:18:09 EST 1997) Minor failures
I was fiddling with formatter until 0955, after that OK. Vc01,2,3 have recycled diodes in 8 and 16MHz bands.

Medicina (Operator \- Mariotti \- at Thu Feb 20 07:30:06 EST 1997) Success
it seems everything OK

Effelsberg (dgraham at Thu Feb 20 08:06:42 EST 1997) Minor failures
I was fiddling with formatter until 0955, after that OK. Vc01,2,3 have recycled diodes in 8 and 16MHz bands.

Jodrell (Alastair Gunn at Thu Feb 20 11:42:07 EST 1997) Success
No known problems. Recorded Mk2 telescope on both MkIV recorder and VLBA recorder. Weather overcast, windy. Tapes NSF00939,885E (MkIV) and EVN00022,12AD (VLBA) sent to Socorro via Fedex 10 minutes after the end of experiment (WOOSH!).

Onsala (John Conway at Fri Feb 21 03:38:26 EST 1997) Failure
High winds, 15m/s, antenna stowed, no observatiosn made, a dummy log will be placed on VLBEER, no tapes will be shipped

Westerbork (Tony FoleyU at Fri Feb 21 08:36:31 EST 1997) Minor failures
We used the single-dish setup with RT6. Linear polarization recorded. Before the session VCs 1-6 had the sideband suppression improved to ~20dB, but the switching diodes were _NOT_ modified. Tape shipped to Socorro on the same day (20Feb) via Fedex.

DSN (Larry Teitelbaum at Mon Feb 24 15:29:14 EST 1997) Success
Goldstone successfully executed the vt9b schedule. The schedule was submitted on time and all scans were recorded. Only one error was logged to the pcfs logfile, a matcn/debug error from the MkIV formatter. The logfile, vt9bg.log, has been copied to the aspen server at Socorro and the tape has been shipped to Socorro. Due to L-band front end problems, our off-source zenith system temperatures are approximately 60-70 K. Also, a noise diode was inadvertently left on during the first three scans which will further raise the system temperature by about 25 K.

Shanghai (Huang Xinyong at Wed Mar 5 04:54:07 EST 1997) Minor failures
Clock offset = 0.08 - 0.09 msec. Formatter slowed down 4 seconds when we change 1 bit sampler to 2 bit sampler accorging to skd file. We adjusted formatter clock after third scan. Weather fine.

Calibration (Fredrik T Rantakyro at Thu May 8 11:34:09 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the Feb97 directory with the name vt9.antab. This file contains the preliminary calibration results and may be updated when I receive updated information from the telescopes and the results from the datareduction of the calibration experiments.